Solve for x: x cubed - 4 cubed = 2x cubed - 7 cubed
Answer by venugopalramana(3286)

x cubed - 4 cubed = 2x cubed - 7 cubed
x^3-4^3=2x^3-7^3
2x^3-7*7*7=x^3-4*4*4
2x^3-x^3=7*7*7-4*4*4=343-64=279
x^3=279
x=cuberoot of 279